OpenStudy (anonymous):

WILL MEDAL AND FAN!! NEED HELP ASAP!! PLEASE!! Describe the graph x3-18x2+101x-180. Find the y and x intercepts

OpenStudy (anonymous):

@mathstudent55 @astrophysics @KEYS

OpenStudy (anonymous):

Please anyone

OpenStudy (anonymous):

@owlcoffee

OpenStudy (anonymous):

Hi!!

OpenStudy (owlcoffee):

\[f:f(x)=x ^{3}-18x^2+101x-180\] Before I bI+ll describe to you why I put the "f:f(x)" there. Very simple, because a graph is but a bunch of points set up in a reference system, where x increases, the value of y will be dependant on the value of "x" that we take in consideration, therefore, the "function" describes the behavior of these points. There are some notable points that most functions have in their composition, which are: (1) The y-intercept (2)The x-interception, often called "zeroes" or "roots". The "y-intercept" is the notable point of the function where the "x" value is zero, therefore, we will take the function and replace all the "x's" with "0's" and that will allow us to know where in the y-axis the point is situated: \[f:f(0)=(0)^3-18(0)^2+101(0)-180\] I'll leave the calculations to you. The "x" interception is a little more complex, but very linear concept, but these notable or notable points have the definition of nullifying the function, meaning that we want to know the values of "x" that make the whole function equal zero: \[x ^{3}-18x^2+101x-180 =0\]

OpenStudy (anonymous):

Is the y intercept (0,-180)

OpenStudy (anonymous):

@owlcoffee

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

@LynFan please help😊

OpenStudy (lynfran):

ok to find the x intercept we let y=0 to find the y intercept we let x=0 and yours y intercept is correct...

OpenStudy (anonymous):

OK so whatvabout the xintercept?

OpenStudy (lynfran):

let y=0 and solve for x ok

OpenStudy (anonymous):

Ok so would it be (-1,0)

OpenStudy (lynfran):

\[x ^{3}-18x ^{2}+101x-180=0\]

OpenStudy (anonymous):

So would there be multiply x intercepts

OpenStudy (anonymous):

Is it 3x+36x-101x-180

OpenStudy (anonymous):

I'm sorry I'm not getting this😥

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

Would the intercepts be (-1,0) (2,0)

OpenStudy (anonymous):

@LynFan yes?

OpenStudy (lynfran):

(5,0), (4,0), (9,0) and (0,-180) the solution to y intercept also... https://www.symbolab.com/solver/algebra-calculator/0%5E%7B3%7D-18%5Cleft(0%5Cright)%5E%7B2%7D%2B101%5Cleft(0%5Cright)-180%3Dy/?origin=button

OpenStudy (anonymous):

Can u showvme how u found that on here? It won't open on my browser

OpenStudy (lynfran):

what r u talking about...how i find what?

OpenStudy (anonymous):

The x intercept

OpenStudy (lynfran):

i type the complete equation here... https://www.symbolab.com/solver/algebra-calculator/x%5E%7B3%7D-18x%5E%7B2%7D%2B101x-180%3D0/?origin=button and the calculator solve for x intercept

OpenStudy (anonymous):

It won't work BC it says I have to pay for it

OpenStudy (anonymous):

Can u just write it here without the link

OpenStudy (lynfran):

ok the first step to do is to factor the equation

OpenStudy (lynfran):

\[x ^{3}-18x ^{2}+101x-180=0\]when factored we get\[(x-4)(x-5)(x-9)=0\] we then let each bracket=0 since they are product of each other..\[(x-4)=0\]\[(x-5)=0\]\[(x-9)=0\]..so \[x=4\]\[x=5\]\[x=9\]

OpenStudy (anonymous):

Then howvwould u graph thaylt

OpenStudy (anonymous):

Can u graph that?

OpenStudy (anonymous):

Wouldn't it u use Descartes sign rule?

OpenStudy (anonymous):

Wouldn't it u use Descartes sign rule?

OpenStudy (anonymous):

Ahhhh that you!!!!

OpenStudy (anonymous):

Thank you!!

OpenStudy (lynfran):

welcome